OpenStudy (anonymous):

Explain the difference between using the sine ratio to solve for a missing angle in a right triangle versus using the cosecant ratio. You must use complete sentences and any evidence needed (such as an example) to prove your point of view. Please help. I'll give you a medal if you help me.

OpenStudy (misssunshinexxoxo):

Do you know what sine and cosecant ratios are?

OpenStudy (anonymous):

yes sin = opp/hyp csc = hyp/opp

OpenStudy (misssunshinexxoxo):

Exactly; an example is csc(30) = 1/sin(30)

OpenStudy (misssunshinexxoxo):

or if it was reversed as well; if you understand the ratios, should be no dilemma to explain it

OpenStudy (anonymous):

Thank you, @misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

My pleasure glad to help; cheers! :)

OpenStudy (triciaal):

yes sin = 1/cosec